OpenStudy (anonymous):

what is lim h tends to 0 hcos(1/h) ?

OpenStudy (anonymous):

This has no definite limit. It goes to infinity.

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

cos(1/0) has any value between -1 to 1 no? then any value between -1 to 1 * 0 should equal to 0 no?

OpenStudy (anonymous):

yes you are right cos(infinity) have value between -1 and 1 so its 0

OpenStudy (anonymous):

so \[\lim h \rightarrow0 hcos1/h \] will be 0 na?

OpenStudy (anonymous):

yup \[\lim(hcos(\frac{ 1 }{ h }))\] when h->0 is 0

OpenStudy (anonymous):

thanks :-)

OpenStudy (anonymous):

no problem